There are four steps in the manufacturing process of a stuffed toy: cutting, stuffing, sealing, and packaging. There are two employees each for cutting and stuffing but one each for sealing and packaging. The processing times of cutting, stuffing, sealing, and packaging are 8, 5, 3, and 2 seconds per toy per employee.

a) What is the cycle time for stage “cutting” (in [min/unit])?

b)What is the cycle time for stage “sealing” (in [min/unit])?

c)Which stage is the bottleneck? What is the process capacity in [units/hour]?

d)If you can add one employee in parallel at the bottleneck stage, what should be the capacity of a new employee so that you increase the process capacity as much as possible?
Answer #1

a) Cycle time for cutting (min/unit)

Time per unit per employee= 8 sec

Number of employees = 2

Thus, 2 units are cut in 8 sec

Cycle time = 4 sec/unit or 4/60 min/unit = 0.067 min/unit

b) Cycle time for sealing (min/unit)

Time per unit per employee= 3 sec

Number of employees = 1

Thus, 1 units is sealed in 3 sec

Cycle time = 3 sec/unit or 3/60 min/unit = 0.05 min/unit

c) Similarly cycle time of stuffing = 5/2/60 min/unit = 0.042 min/unit

cycle time of packaging = 2/60 = 0.033 min/unit

Thus, bottleneck us Cutting process with cycle time of 0.067 min/unit or 4 sec/unit

Process capacity in an hour = 60*60/4 = 900 units per hour

d) Cutting requires 4 sec per unit as of now. Sealing requires 3 sec per unit

Thus, new employee should be able to bring cutting output to 3 sec per unit. So 8/(2+x) = 3

x = 2/3 sec for 1 unit

x = 12 sec per unit

So, new employee should be able to do cutting in 12 sec per unit